Solve for x
8/x=28/42
: Solve for x
8/x=28/42

Answer by jim_thompson5910(9226)


8/x=28/42 Start with the given ratio


8=(28/42)*x Multiply both sides by x


8*42=28*x Multiply both sides by 42


336=28*x Multiply


336/28=x Divide both sides by 28


12=x Simplify


So our answer is x=12